### v1

@@ B.3 Rotationally Invariant Encoding for $S_{4}$
Figure 14: The action $(U\otimes U\otimes U\otimes U)$ written in the Schur basis kirby_thesis ; Kirby_2018 . Blocks correspond to irreducible subspaces of $SU(2)^{\otimes 4}$. The dimension of a block is the dimension of the irreducible subspace and the amount of repeats is the multiplicity. The trivial representation of $SU(2)^{\otimes 4}$ corresponds to blocks of $[1]$, with multiplicity $2$. This corresponds to a $2$-dimensional subspace in $S_{n}$ with multiplicity $1$.
Using the work of kirby_thesis ; Kirby_2018 we consider the group $S_{4}$. In this case the action of $U\otimes U\otimes U\otimes U$ when written in the Schur basis is shown in Figure 14. Within this decomposition, we can identify that there are two basis elements in the Schur basis that are invariant under the action of $(U\otimes U\otimes U\otimes U)$. This is the trivial representation of $SU(2)^{\otimes 4}$ which has dimension equal to 1, but has a multiplicity equal to 2. This correspond to the basis states
@@
Due to the Schur-Weyl duality, we know that one of the irreducible subspaces of $S_{n}$ will correspond to the same basis states. However, in this case the multiplicity of the irreducible representation of $S_{n}$ will be 1 and the dimension will equal 2. One can find that by projecting onto the $r=3$ representation of $S_{4}$ (as specified in Table 2) using the projection circuit proposed in Section 4, the quantum input state $\lvert\psi\rangle$ will projected onto some state $\lvert\psi_{\text{rot}}\rangle=\lambda_{1}\lvert d_{1}\rangle+\lambda_{2}\lvert d_{2}\rangle$. In this case there are now two basis states that can be used. However, due to the normalisation constraint $\rvert\lambda_{1}\rvert^{2}+\rvert\lambda_{2}\rvert^{2}=1$ the effective dimension is reduced to be equal to 1 overall. This means that at least some information from the original input state is retained when the projection is performed and is therefore a non-trivial example, although effective projection to a $1$-dimensional space may be too great a dimensionality reduction in practice.
In general, if it is possible to calculate the multiplicity of the trivial representation of dimension 1 of $SU(2)^{\otimes n}$ to be $m_{SU}$, then by Schur-Weyl duality there exists an irreducible subspace of $S_{n}$ of dimension $m_{SU}$ and multiplicity of $1$. Projecting to this subspace will therefore result in reducing the dimensionality of the encoded data to $m_{SU}-1$, after considering the normalisation condition. We provide this rotationally invariant encoding as an example to demonstrate how the irreducible subspace projection circuit could be utilised for point cloud data, complementing previous work on permutation invariant encodings of point cloud data heredge2023permutation . It may be the case that this is not feasible or useful for $n>4$ and such investigations could be subject to further research. However, the main motivation was to provide an additional example of how the projection framework could be used to encode symmetry in a model. In general, the irreducible subspace projection framework we introduced can be used for any representation of any finite compact group $G$ along with any encoding procedure for $\lvert\psi\rangle$ and hence there exists ample flexibility for application to a wide array of problems beyond those mentioned as examples in this work.
@@ Appendix C Point Cloud Numeric Details
This section specifies the details of the numerical results reported in Section 4.4. To generate the data, $n=3$ points were sampled from the surface of a shape, either a sphere or a torus as originally used in heredge2023permutation , to create point clouds. The process is repeated with an equal amount of point clouds from each shape until the desired number of total samples is reached. These are split into training sets (80%) and testing sets (20%). We evaluated the performance of the model on the test set and averaged the accuracy in 10 experiments in which a newly generated dataset is used each time. Both the sphere and torus are centred at the origin, with the torus scaled to match the sphere’s average point magnitude. All data is normalised between $-\frac{\pi}{2}$ and $\frac{\pi}{2}$ to be encoded as rotation angles.
To encode a point cloud, each point $\mathbf{p}_{i}=(x_{i},y_{i},z_{i})$ is first encoded into a two qubit quantum state $\lvert\mathbf{p}_{i}\rangle$ using two repeated layers of the encoding circuit shown in Figure 15. Then the corresponding statevector is found using Qiskit statevector_simulator qiskit . From here we can write the entire point cloud as $\lvert P\rangle=\lvert\mathbf{p}_{1}\rangle\otimes\lvert\mathbf{p}_{2}\rangle\otimes\lvert\mathbf{p}_{3}\rangle$. In order to perform the simulations we then found the projection of $\lvert P\rangle$ onto each irreducible subspace classically utilising the result specified in Theorem 3 and the character table for $S_{3}$. From here the $a_{r}$ parameters specified in Equation 110 are used to produce the correct linear combination of projections and arrive at the partially symmetric subspace amplified state $\lvert P\rangle_{\alpha}$, with the amount of symmetry parameterised by $\alpha$. We then implement a quantum support vector machine classification by calculating the inner product between the data points to form a kernel matrix, which is then input to a classical support vector machine for the final classification havlicek_supervised_2019 ; heredge2023permutation . We recorded the average test accuracy over 10 experiments for a given $\alpha$ and then repeated with different $\alpha$ values and plotted the results to see which value of $\alpha$ is optimal for the data.
Figure 15: One layer of the encoding circuit used to encode an individual point $\mathbf{p}_{i}=(x_{i},y_{i},z_{i})$ in the symmetric subspace amplification model. The function $\phi(x_{i},y_{i},z_{i})=\frac{2}{\pi^{2}}(\pi-x_{i})(\pi-y_{i})(\pi-z_{i})$ means the entanglement of the state is dependent on some polynomial of all input variables. To generate the encoded state $\lvert\mathbf{p}_{i}\rangle$ two repeated layers of this circuit are used havlicek_supervised_2019 ; Heredge21 .
@@ Appendix D Dimensionality Reduction in Pooling Layers
Regarding the average pooling layers implemented in this work in Section 3, it is crucial to understand that we have confined our operations to averaging quantum states across image translations. Notably, binary addition operates cyclically, where adding 1 to the highest value loops back to the lowest value, meaning our tiling would wrap around the edges of the image to perform the averaging. This may be an undesirable effect, unless the image has some periodic feature. It is also the case that the average pooling described does not provide any dimensionality reduction, as we can calculate the pooling window average for every pixel at no additional computational cost.
@@
Clearly as this is a probabilistic procedure we expect that significantly more efficient implementations are possible depending on the exact specifics of the problem. For example, if it is known that some of the most significant qubits correspond to pixels that should not be used, then they can simply be discarded without needing to implement flag operators and extra ancilla qubits. Implementation of the discarding part of the pooling process will therefore be dependent on the exact requirements of the model, although in the worst-case scenario resorting to using $F_{\text{FLAG}}$ operators, as described above, to remove specific states would succeed in reducing the image dimensionality.
Appendix E Input Skip Connections Only ResNet
@@
$P_{\text{PREP}}=\sum_{f=1}^{L}\gamma_{f}\lvert b_{f}\rangle\langle b_{1}\rvert+\sum_{j=2}^{L}\sum_{f=1}^{L}\delta_{jf}\lvert b_{f}\rangle\langle b_{j}\rvert,$ (142)
@@
$=\sum_{f=1}^{L}\gamma_{f}\lvert b_{f}\rangle\Big{(}\prod_{l=f}^{M}W_{l}(\mathbf{\theta}_{l})\lvert\phi(\mathbf{x})\rangle\Big{)}.$ (145)
@@
$(P_{\text{PREP}})^{\dagger}=\sum_{f=1}^{L}\gamma_{f}^{*}\lvert b_{1}\rangle\langle b_{f}\rvert+\sum_{j=2}^{L}\sum_{f=1}^{L}\delta_{jf}^{*}\lvert b_{j}\rangle\langle b_{f}\rvert,$ (146)
@@
$=\sum_{f=1}^{L}\prod_{l=f}^{L}\rvert\gamma_{f}\rvert^{2}\lvert b_{1}\rangle W_{l}(\mathbf{\theta}_{l})\lvert\phi(\mathbf{x})\rangle+\sum_{j=2}^{L}\lvert b_{j}\rangle(...).$ (147)
@@
$\pi_{S}=\rvert\sum_{f=1}^{L}\prod_{l=f}^{L}\rvert\gamma_{f}\rvert^{2}W_{l}(\mathbf{\theta}_{l})\lvert\phi(\mathbf{x})\rangle\rvert^{2},$ (148)
@@
$=\frac{1}{\Omega^{\prime}}\sum_{f=1}^{L}\prod_{l=f}^{L}\rvert\gamma_{f}\rvert^{2}W_{l}(\mathbf{\theta}_{l})\lvert\phi(\mathbf{x})\rangle,$ (149)
@@
This LCU method will successfully prepare the desired residual network. Hence we have shown that VQC models can be built with skipped connections under the Quantum ResNet framework by utilising the LCU method to implement these non-unitary operations. An example circuit implementation for the four-layer case $L=4$ is shown in Figure 16. Notably, this method is more qubit efficient requiring $\mathcal{O}(\log(L))$ qubits for $L$ layers, as opposed to $\mathcal{O}(L)$ qubits in the previous case. It also results in a quadratic (in $L$) number of terms in the final loss function, which means that the absence of barren plateau argument made in the main text can hold as long as $L=\mathcal{O}(\text{poly}(n))$. This is in contrast to the framework in the main text which in general could contain an exponential (in $L$) amount of terms in the final loss function. Due to the fact that this could exponentially suppress any beneficially scaling terms, the main text implementation would require that $L=\mathcal{O}(\log(n))$ in order to avoid this.
@@
$\pi_{S}=$ $\rvert\sum_{f=1}^{L}\prod_{l=f}^{L}\rvert\gamma_{f}\rvert^{2}W_{l}(\mathbf{\theta}_{l})\lvert\phi(\mathbf{x})\rangle\rvert^{2}$
@@
E.1 Two Gate Input Worked Example

### v2

@@ B.3 Rotationally Invariant Encoding for $S_{4}$
Figure 21: The action $(U\otimes U\otimes U\otimes U)$ written in the Schur basis Kirby ; Kirby and Strauch (2018). Blocks correspond to irreducible subspaces of $SU(2)^{\otimes 4}$. The dimension of a block is the dimension of the irreducible subspace and the amount of repeats is the multiplicity. The trivial representation of $SU(2)^{\otimes 4}$ corresponds to blocks of $[1]$, with multiplicity $2$. This corresponds to a $2$-dimensional subspace in $S_{n}$ with multiplicity $1$.
Using the work of Kirby ; Kirby and Strauch (2018) we consider the group $S_{4}$. In this case the action of $U\otimes U\otimes U\otimes U$ when written in the Schur basis is shown in Figure 21. Within this decomposition, we can identify that there are two basis elements in the Schur basis that are invariant under the action of $(U\otimes U\otimes U\otimes U)$. This is the trivial representation of $SU(2)^{\otimes 4}$ which has dimension equal to 1, but has a multiplicity equal to 2. This correspond to the basis states
@@
Due to the Schur-Weyl duality, we know that one of the irreducible subspaces of $S_{n}$ will correspond to the same basis states. However, in this case the multiplicity of the irreducible representation of $S_{n}$ will be 1 and the dimension will equal 2. One can find that by projecting onto the $r=3$ representation of $S_{4}$ (as specified in Table 2) using the projection circuit proposed in Section IV, the quantum input state $\lvert\psi\rangle$ will projected onto some state $\lvert\psi_{\text{rot}}\rangle=\lambda_{1}\lvert d_{1}\rangle+\lambda_{2}\lvert d_{2}\rangle$. In this case there are now two basis states that can be used. However, due to the normalisation constraint $\rvert\lambda_{1}\rvert^{2}+\rvert\lambda_{2}\rvert^{2}=1$ the effective dimension is reduced to be equal to 1 overall. This means that at least some information from the original input state is retained when the projection is performed and is therefore a non-trivial example, although effective projection to a $1$-dimensional space may be too great a dimensionality reduction in practice.
In general, if it is possible to calculate the multiplicity of the trivial representation of dimension 1 of $SU(2)^{\otimes n}$ to be $m_{SU}$, then by Schur-Weyl duality there exists an irreducible subspace of $S_{n}$ of dimension $m_{SU}$ and multiplicity of $1$. Projecting to this subspace will therefore result in reducing the dimensionality of the encoded data to $m_{SU}-1$, after considering the normalisation condition. We provide this rotationally invariant encoding as an example to demonstrate how the irreducible subspace projection circuit could be utilised for point cloud data, complementing previous work on permutation invariant encodings of point cloud data Heredge et al. (2024). It may be the case that this is not feasible or useful for $n>4$ and such investigations could be subject to further research. However, the main motivation was to provide an additional example of how the projection framework could be used to encode symmetry in a model. In general, the irreducible subspace projection framework we introduced can be used for any representation of any finite group $G$ along with any encoding procedure for $\lvert\psi\rangle$ and hence there exists ample flexibility for application to a wide array of problems beyond those mentioned as examples in this work.
@@ Appendix C Point Cloud Numeric Details
This section specifies the details of the numerical results reported in Section IV.4. To generate the data, $n=3$ points were sampled from the surface of a shape, either a sphere or a torus as originally used in Heredge et al. (2024), to create point clouds. The process is repeated with an equal amount of point clouds from each shape until the desired number of total samples is reached. These are split into training sets (80%) and testing sets (20%). We evaluated the performance of the model on the test set and averaged the accuracy in 10 experiments in which a newly generated dataset is used each time. Both the sphere and torus are centred at the origin, with the torus scaled to match the sphere’s average point magnitude. All data is normalised between $-\frac{\pi}{2}$ and $\frac{\pi}{2}$ to be encoded as rotation angles.
To encode a point cloud, each point $\mathbf{p}_{i}=(x_{i},y_{i},z_{i})$ is first encoded into a two qubit quantum state $\lvert\mathbf{p}_{i}\rangle$ using two repeated layers of the encoding circuit shown in Figure 22. Then the corresponding statevector is found using Qiskit statevector_simulator Abraham et al. (2019). From here we can write the entire point cloud as $\lvert P\rangle=\lvert\mathbf{p}_{1}\rangle\otimes\lvert\mathbf{p}_{2}\rangle\otimes\lvert\mathbf{p}_{3}\rangle$. In order to perform the simulations we then found the projection of $\lvert P\rangle$ onto each irreducible subspace classically utilising the result specified in Theorem 3 and the character table for $S_{3}$. From here the $a_{r}$ parameters specified in Equation 115 are used to produce the correct linear combination of projections and arrive at the partially symmetric subspace amplified state $\lvert P\rangle_{\alpha}$, with the amount of symmetry parameterised by $\alpha$. We then implement a quantum support vector machine classification by calculating the inner product between the data points to form a kernel matrix, which is then input to a classical support vector machine for the final classification Havlíček et al. (2019); Heredge et al. (2024). We recorded the average test accuracy over 10 experiments for a given $\alpha$ and then repeated with different $\alpha$ values and plotted the results to see which value of $\alpha$ is optimal for the data.
Figure 22: One layer of the encoding circuit used to encode an individual point $\mathbf{p}_{i}=(x_{i},y_{i},z_{i})$ in the symmetric subspace amplification model. The function $\phi(x_{i},y_{i},z_{i})=\frac{2}{\pi^{2}}(\pi-x_{i})(\pi-y_{i})(\pi-z_{i})$ means the entanglement of the state is dependent on some polynomial of all input variables. To generate the encoded state $\lvert\mathbf{p}_{i}\rangle$ two repeated layers of this circuit are used Havlíček et al. (2019); Heredge et al. (2021).
@@ Appendix D Dimensionality Reduction in Pooling Layers
Regarding the average pooling layers implemented in this work in Section III, it is crucial to understand that we have confined our operations to averaging quantum states across image translations. Notably, binary addition operates cyclically, where adding 1 to the highest value loops back to the lowest value, meaning our tiling would wrap around the edges of the image to perform the averaging. This may be an undesirable effect, unless the image has some periodic feature. It is also the case that the average pooling described does not provide any dimensionality reduction, as we can calculate the pooling window average for every pixel at no additional computational cost.
@@
Clearly as this is a probabilistic procedure we expect that significantly more efficient implementations are possible depending on the exact specifics of the problem. For example, if it is known that some of the most significant qubits correspond to pixels that should not be used, then they can simply be discarded without needing to implement flag operators and extra ancilla qubits. Implementation of the discarding part of the pooling process will therefore be dependent on the exact requirements of the model, although in the worst-case scenario resorting to using $F_{\text{FLAG}}$ operators, as described above, to remove specific states would succeed in reducing the image dimensionality. We refer the reader to Wei et al. (2022) for more details on implementing pooling within a full quantum convolutional neural network framework.
Appendix E Subtraction Circuit Scaling
The subtraction / decrement operator we consider is controlled by a single ancilla qubit meaning that overall we consider $n+1$ qubits consisting of a total of $n$ multi-controlled Toffolli gates, with the largest being controlled by $n$ qubits, and the amount of controlling qubits decreasing by $1$ each time such that the final gate is a CNOT gate. For $n+1$ total qubits the $n$ qubit multi-controlled Toffoli can be decomposed into at worst $\mathcal{O}(n^{2})$ basic operations Barenco et al. (1995), which may be improved if ancillas are used Baker et al. (2019). The $m$ qubit controlled Toffoli with $m<n-1$ can be created by $\mathcal{O}(m)$ basic operations Barenco et al. (1995). As there are $n-1$ gates of this type in total, their overall contribution will be bounded by $\mathcal{O}(n^{2})$. Hence, overall we can upper bound the required number of basic operations at $\mathcal{O}(n^{2})$ to create the basic controlled subtraction gate (Although it has also been suggested linear scaling may be possible for increment/decrement gates using extra ancillas Gidney ; Khattar and Gidney (2024); Gidney (2018)). Note that as we consider subtraction operators where the subtraction amount doubles each time, this means subsequent subtraction operators can be implemented by ignoring an additional least significant qubit as the subtraction amount doubles. This means that in practise subsequent subtraction operators will use even fewer gates, although we will not include this in our complexity calculation. As $n=\log(N)$ and as we require $\mathcal{O}(\log(D))$ operators in total then overall we require $\mathcal{O}(\log(D)(\log(N))^{2})$ basic operations to implement the average pooling.
Appendix F Degeneracy Avoidance in Average Pooling
If window length $D<2^{L}$ where $L$ is the number of ancilla qubits and controlled operators then we need to change the final subtraction operator value. For example if $D=8$ then we would we would generate all $8$ operators $\mathbb{I}+\hat{T}_{1}+\hat{T}_{2}+...+\hat{T}_{7}$ by using only three controlled unitaries $\hat{T}_{1},\hat{T}_{2},\hat{T}_{3}$. However if we had $D=7$ then we could attempt to implement this using $\hat{T}_{1},\hat{T}_{2},\hat{T}_{3}$ but writing this out in full one will see this implements
$\mathbb{I}+\hat{T}_{1}+\hat{T}_{2}+2\hat{T}_{3}+\hat{T}_{4}+\hat{T}_{5}+\hat{T}_{6}$ (150)
which correctly has only $D=7$ terms, but there is a degeneracy for $\hat{T}_{3}$ meaning that it has double weighting and we do not have an equal superposition. This may however be preferable if one wishes to create a general convolutional filter, where in this case the pixel corresponding to $\hat{T}_{3}$ in the convolutional filter is supposed to be given a weighting double that of all others.
If one wishes to avoid this degeneracy though, then it can be achieved by avoiding an equal superposition of all ancilla qubits and instead initialising them to avoid repeating terms. In the above example both the ancilla states $\lvert 100\rangle$ and $\lvert 110\rangle$ result in an implementation of $\hat{T}_{3}$. Setting the amplitude of either of these states to zero would remove the degeneracy and correctly implement the average pooling layer again.
Appendix G Input Skip Connections Only ResNet
@@
$P_{\text{PREP}}=\sum_{f=1}^{L}\gamma_{f}\lvert b_{f}\rangle\langle b_{1}\rvert+\sum_{j=2}^{L}\sum_{f=1}^{L}u_{j,f}\lvert b_{f}\rangle\langle b_{j}\rvert,$ (151)
@@
$=\sum_{f=1}^{L}\gamma_{f}\lvert b_{f}\rangle\Big{(}\prod_{l=f}^{L}W_{l}(\mathbf{\theta}_{l})\lvert\phi(\mathbf{x})\rangle\Big{)}.$ (154)
@@
$(P_{\text{PREP}})^{\dagger}=\sum_{f=1}^{L}\gamma_{f}^{*}\lvert b_{1}\rangle\langle b_{f}\rvert+\sum_{j=2}^{L}\sum_{f=1}^{L}u_{jf}^{*}\lvert b_{j}\rangle\langle b_{f}\rvert,$ (155)
@@
$=\sum_{f=1}^{L}\rvert\gamma_{f}\rvert^{2}\lvert b_{1}\rangle\prod_{l=f}^{L}W_{l}(\mathbf{\theta}_{l})\lvert\phi(\mathbf{x})\rangle+\sum_{j=2}^{L}\lvert b_{j}\rangle(...).$ (156)
@@
$\pi_{S}=\rvert\sum_{f=1}^{L}\rvert\gamma_{f}\rvert^{2}\prod_{l=f}^{L}W_{l}(\mathbf{\theta}_{l})\lvert\phi(\mathbf{x})\rangle\rvert^{2},$ (157)
@@
$=\frac{1}{\Omega^{\prime}}\sum_{f=1}^{L}\rvert\gamma_{f}\rvert^{2}\prod_{l=f}^{L}W_{l}(\mathbf{\theta}_{l})\lvert\phi(\mathbf{x})\rangle,$ (158)
@@
This LCU method will successfully prepare the desired residual network. Hence we have shown that VQC models can be built with skipped connections under the quantum ResNet framework by utilising the LCU method to implement these non-unitary operations. An example circuit implementation for the four-layer case $L=4$ is shown in Figure 23. Notably, this method is more qubit efficient requiring $\mathcal{O}(\log(L))$ qubits for $L$ layers, as opposed to $\mathcal{O}(L)$ qubits in the previous case.
@@
$\pi_{S}=$ $\rvert\sum_{f=1}^{L}\rvert\gamma_{f}\rvert^{2}\prod_{l=f}^{L}W_{l}(\mathbf{\theta}_{l})\lvert\phi(\mathbf{x})\rangle\rvert^{2}$
@@
G.1 Two Gate Input Worked Example